OpenStudy (smurfy14):

2log(base8)4-(1/3)log(base8)8=x

OpenStudy (anonymous):

What number are both 8 and 4 a power of?

myininaya (myininaya):

log(base8)4^2-log(base8)8^(1/3)=log(base8)[4^2/8^(1/3)]

OpenStudy (anonymous):

Are you good now smurfy?

OpenStudy (smurfy14):

sorry for late response i didnt see my new notifications! um i get what myininaya did but what would you do after that?

myininaya (myininaya):

simplify!

myininaya (myininaya):

=log(base8)[16/2] i will let you do the last step

myininaya (myininaya):

or last two steps

OpenStudy (smurfy14):

so x=1 correct?

myininaya (myininaya):

:) YES!

OpenStudy (smurfy14):

whoo thanks so much!

OpenStudy (smurfy14):

think you could help me with one more?

myininaya (myininaya):

ok

OpenStudy (smurfy14):

2. Jamal has a game with 2 groups of tiles. The first group of 26 tiles is labeled with all the letter of the alphabet, The second group of 10 tiles is #0-9. If Jamal draws one letter tile and one # tile at the random, what is the probability that he will draw a letter in his name and an odd number? A. (1/13) B. (5/52) C. (1/4) D. (7/26)

myininaya (myininaya):

i got 17/26 :(

myininaya (myininaya):

let me think about it

OpenStudy (smurfy14):

ok and ya all the answers tht i was getting werent on there either :P

myininaya (myininaya):

i suck at probability, but give me a few and I will think

OpenStudy (anonymous):

I get A

OpenStudy (smurfy14):

could u plz explain it to me?

myininaya (myininaya):

did you P(AUB)=P(A)+P(B)-P(A and B)

OpenStudy (anonymous):

No, I used the probability of two independent events is their product.

OpenStudy (anonymous):

So how many ways can Jamal draw a letter that is in his name?

OpenStudy (smurfy14):

jaml

myininaya (myininaya):

its independent events because A has nothing to do with B good thinking bond

OpenStudy (anonymous):

right, so 4/26

OpenStudy (smurfy14):

someone explain how rhey got A!

myininaya (myininaya):

dont worry about A

OpenStudy (anonymous):

Now, how about the probability of pulling an odd tile?

OpenStudy (smurfy14):

1/10

myininaya (myininaya):

how many odd numbers are between 0 and 9?

OpenStudy (smurfy14):

ohhhh ODD 5/10!

myininaya (myininaya):

ok so since they are independent events you multiply the (5/10)(4/26)

OpenStudy (smurfy14):

ok see ididnt get it bc ithought it said 1/10 not 5/10 imissedthe odd part lol thanks!!

myininaya (myininaya):

lol

OpenStudy (smurfy14):

one more q! 3^2x=7^(x-1)

OpenStudy (anonymous):

Eeek, that doesn't look good...

OpenStudy (smurfy14):

hahaha

myininaya (myininaya):

q! is read as q factorial?

myininaya (myininaya):

jk lol i know you mean another question!

myininaya (myininaya):

take the natural log of both sides

OpenStudy (smurfy14):

you mean get the decimal form of both? if so thats what i did but for some reason i came out with a negative # :(

myininaya (myininaya):

ln3^(2x)=ln7^(x-1)

myininaya (myininaya):

2xln3=(x-1)ln7

myininaya (myininaya):

2xln3=xln7-ln7

myininaya (myininaya):

2xln3-xln7=-ln7

myininaya (myininaya):

x(2ln3-ln7)=-ln7

myininaya (myininaya):

x=-ln7/(2ln3-ln7)

myininaya (myininaya):

x=-ln7/(ln8-ln7)

myininaya (myininaya):

x=-ln7/(ln(8/7))

OpenStudy (smurfy14):

ok i see so b4 i write the work does the answer comeout positive?

myininaya (myininaya):

you have a negative in front of it so its negative

OpenStudy (smurfy14):

ok so were allowed to have a negative as the answer?

myininaya (myininaya):

ln(negative)=does not exist ln(0) does not exist -lna exist if a>0

OpenStudy (smurfy14):

so the answer is?

myininaya (myininaya):

the answer to?

myininaya (myininaya):

are we still talking about the previous solution?

OpenStudy (smurfy14):

ya the one with the negative

myininaya (myininaya):

so you are saying its not a solution? why do you say that?

OpenStudy (smurfy14):

what? i just want to know what you got for the answer lol

myininaya (myininaya):

we can check it to see

OpenStudy (smurfy14):

youre confusing me

myininaya (myininaya):

you want to know if it is a solution. if you ever have doubt about something, check it

OpenStudy (smurfy14):

no i dont want to know if it is a solution i just want to know the answer lol so would the answer be "no solution"?

myininaya (myininaya):

give me just sec and i will check k??

myininaya (myininaya):

it is a solution if i did each step correctly

OpenStudy (smurfy14):

so the final answer is....?

OpenStudy (smurfy14):

you know like x=?

myininaya (myininaya):

i can scroll back up and look just sec

myininaya (myininaya):

it says x=-ln(7)/(ln(8/7))

OpenStudy (smurfy14):

so if you put that in your calc what is the number? (sorry i dont have a good calc with me)

myininaya (myininaya):

> evalf(solve(3^(2*x) = 7^(x-1))); -7.742930467 >

OpenStudy (smurfy14):

ok thanks!

myininaya (myininaya):

smurfy it is not easy to check the solution, but it is the solution because exponential functions are continuous everywhere

myininaya (myininaya):

but logarithmic functions are not continuous everywhere log(baseb)x=lnx/lnb so whenever the base is 1 or negative or 0 and whenever x is negative or 0, the the function is undefined